# AutoCV: Empowering Reasoning with Automated Process Labeling via Confidence Variation

### v2

@@ Abstract
In this work, we propose a novel method named Automated Process Labeling via Confidence Variation (AutoCV) to enhance the reasoning capabilities of large language models (LLMs) by automatically annotating the reasoning steps. Our approach begins by training a verification model on the correctness of final answers, enabling it to generate automatic process annotations. This verification model assigns a confidence score to each reasoning step, indicating the probability of arriving at the correct final answer from that point onward. We detect relative changes in the verification’s confidence scores across reasoning steps to automatically annotate the reasoning process. This alleviates the need for numerous manual annotations or the high computational costs associated with model-induced annotation approaches. We experimentally validate that the confidence variations learned by the verification model trained on the final answer correctness can effectively identify errors in the reasoning steps. Subsequently, we demonstrate that the process annotations generated by AutoCV can remarkably enhance the accuracy of the verification model in selecting the correct answer from multiple outputs generated by LLMs. Notably, we achieve substantial improvements across five datasets in mathematics and commonsense reasoning. The source code of AutoCV is available at https://github.com/rookie-joe/AUTOCV.
[…]
@@ 1 Introduction
Large language models (LLMs) have shown impressive performance on various reasoning tasks [1, 2, 3, 4]. Prior efforts primarily focus on specific prompting techniques, such as few-shot prompting with intermediate steps and augmented demonstrations [5, 6, 7, 8]. While these methods have shown promise, their effectiveness is often task-specific, and designing prompts can be labor-intensive, leading to inconsistent results [9, 10]. Another approach to improve reasoning in LLMs is through instruction tuning or knowledge distillation [11, 12, 13, 14]. These methods typically involve fine-tuning LLMs and require a large set of examples annotated with chain-of-thoughts (CoT). However, these approaches can be resource-intensive and may not always produce reliable results.
@@
To address these challenges, verification techniques have emerged as a promising solution [15, 16]. Verification models are trained to evaluate and potentially correct the reasoning process generated by LLMs. This approach aims to mitigate the risk of relying solely on the top-1 result, which may not always be reliable [17, 18]. By reranking candidate responses, verification models can ensure higher accuracy and consistency in LLM outputs. Moreover, they provide valuable feedback for improving LLMs [19, 20] further.
@@
Verification models generally fall into two training paradigms: outcome supervision and process supervision. In outcome supervision, the training annotations rely on the correctness of the final answer [21, 22], while in process supervision, the annotations are based on evaluations of each reasoning step [23, 19]. However, process supervision is demanding in terms of annotations. Typically, it relies on either expensive and highly skilled human evaluators [23, 16] or model-induced process annotations [18, 17] to estimate the future correctness of the current reasoning step using Monte Carlo tree search [24, 25]. In contrast, outcome supervision only requires annotations for the output, making it more economical in terms of annotation effort but less effective. That being said when answers involve multiple reasoning paths, aforementioned methods require numerous samples to ensure accurate estimations.
@@
In this paper, we present a novel method named Automated Process Labeling via Confidence Variation (AutoCV), which enjoys both advantages of process supervision and output supervision. Our method starts by utilizing outcome supervision annotations to learn an outcome-supervised verification model. Then, this model assigns step-level scores to unannotated intermediate reasoning steps, which estimate the confidence in each reasoning step’s likelihood of leading to a correct final answer. By calculating the relative confidence variation, we could produce process annotations with significantly reduced annotation effort. With our process supervision data, we train an LLM model, resulting in a process-supervised enhanced verification model. The overall framework of our method is illustrated in Figure 1. We conduct extensive experiments across five datasets, including mathematical reasoning benchmarks and commonsense reasoning tasks. The results demonstrate that our method effectively improves the reasoning capability of the model with our highly efficient labeling scheme for process supervision, thus accomplishing two tasks with a single strategy. In summary, our contribution is as follows:
  • We introduce AutoCV, a method for automatically labeling process data to enhance the reasoning capabilities of LLMs. Our approach combines the strengths of output supervision and process supervision to annotate reasoning steps automatically.
@@
  • AutoCV effectively identifies variations in model confidence to annotate the correctness of intermediate reasoning steps, enabling efficient automatic labeling for process supervision.
@@
  • Comprehensive experiments demonstrate that process supervision data generated by AutoCV significantly improves the performance and scalability of verification models in mathematical and commonsense reasoning tasks, while greatly reducing the need for manual intervention and extensive computational resources.
@@
Figure 1: An overview of our framework. Our approach utilizes an outcome-supervised verifier to automatically generate process annotations for each reasoning step by detecting its own confidence variations, without relying on ground truth annotations. Our method efficiently produces annotations serving as process supervision during the LLM training, which sidesteps costly annotations.
@@ 2 Related Works
To enhance the reasoning capabilities of LLMs, prior research primarily focuses on specific prompting techniques [26]. Existing efforts include few-shot prompting with intermediate steps augmented demonstrations [5, 8, 27] or zero-shot prompting with specific instructions [28, 29]. Although these methods have shown promising results, their effectiveness is often constrained by their task-specific nature and the labour-intensive process of designing prompts, leading to inconsistent outcomes across different tasks [9, 10]. Another strategy to facilitate reasoning involves instruction tuning or knowledge distillation, which elicits reasoning paths from LLMs without explicit prompting [12, 13, 30, 31]. These approaches typically involve resource-intensive fine-tuning over LLMs and require a large set of examples annotated with chain-of-thoughts (CoT). Unlike methods that directly modify parameters or prompts, AutoCV focuses on training an additional verification model to select the desired output from the original model’s output. This approach is further discussed in the context of process supervision in the following paragraph.
@@
Recent efforts have focused on enhancing the reasoning capabilities of LLMs through the use of verifiers to select the best answer from multiple candidates. There are two main types of verifiers: the Outcome-Supervised Verifier (OSV) and the Process-Supervised Verifier (PSV). The OSV is supervised with a signal based on the final answer [21, 22], while the PSV is with detailed feedback which requires evaluating individual reasoning steps [15, 19, 16, 23]. Despite the time-consuming annotation cost, the PSV offers several advantages that make it preferable to the OSV. It can provide fine-grained feedback by pinpointing the location of errors, which is valuable for reinforcement learning and automatic correction [16, 20]. To alleviate the extensive human annotation, recent approaches [17, 18] propose a machine annotation framework using Monte Carlo Tree Search [24, 25]. This annotation process demands a lot of computing resources, potentially imposing a limitation on the usage. AutoCV is more efficient, as it utilizes an outcome-supervised verification model to assign confidence to each reasoning step and calculate relative confidence variations, eliminating the need for additional sampling or manual labeling.
3 Method
In the following, we first introduce the problem setting in Section 3.1. Then we discuss the motivation behind why we believe it is necessary to train a verification model in Section 3.2. Finally, we describe how we transition from outcome supervision to process supervision during the training of the verification model in Section 3.3.
@@ 3.1 Problem Setting
Given an LLM as a response generator to produce multiple responses for a given question, we aim to have an effective method to help us select a correct response from the multiple candidates. In other words, our goal is to maximize the probability of choosing a correct solution from the multiple options. We refer to this as response selection.
We also outlier the notations utilized in our following training framework. $q$: The specific question posed to the model. $S^{(1:t)}_{i}$: The intermidiate reasoning steps for question $q$ up to and including step $t$ within the $i$-th solution. $y_{i}$: The label of correctness for the $i$-th solution. $\mathcal{Q}$: The dataset of questions used for training.
@@ 3.2 Motivation
The methods for response selection can be categorized into two main types: models fine-tuned specifically for the choice task and those utilizing different prompting strategies.
In our exploration, we initially investigate whether existing open-source LLMs could serve as effective selection agents to evaluate model outputs and choose the correct response without fine-tuning. We choose Mixtral-Instruct [32] as the response generator. We employ the metric pass@k to evaluate model performance, defined as the scenario where at least one correct instance is within the model’s first k attempts. The evaluation results of the response from Mixtral-Instruct are detailed in Table 1. To arrive at a comprehensive and reliable conclusion, we then employ models of different sizes as the selector, ranging from 7b to over 70b parameters. And we apply different prompting strategies to assess their performance on the task of response selection. The experiment results, summarized in Table 2, showcase the comparison of different prompt strategies across different model sizes tested on GSM8K test set [21]:
Table 1: Performance of Mixtral-Instruct on the GSM8K test set. All results are reported in %.
@@
Table 2: Comparison of different selection methods across various model sizes (from 7B to 72B) for selecting a response from candidate responses generated by Mixtral-Instruct. All results are reported in %.
@@
It can be observed that without fine-tuning and relying solely on prompting, even models with over 70 billion parameters cannot yield efficient performance in the response selection task. In the following discussion, we will introduce our approach to enhance the performance of LLMs in response selection. To this end, we will concentrate on the training and improvement of verification models for this specific task.
@@ 3.3 Training Methodology
Our proposed AutoCV training approach combines outcome supervision with process supervision effectively. We describe each step of AutoCV in detail below.
Outcome-Supervision Initially, we train an outcome-supervised verification model based on whether the final answer in the solution is correct or not, following [22]. We denote the outcome-supervised verification (OSV) model as $f_{\bm{\theta}}(\cdot)$, where $\bm{\theta}$ represents the optimized parameters. The Mean Squared Error (MSE) loss used for outcome supervision for each solution step is defined as follows:
@@
The overall objective function, incorporating the entire set of training questions $\mathcal{Q}$, is expressed as:
@@
Where $n$ denotes the total number of solutions proposed for each question, and $m_{i}$ denotes the total number of reasoning steps within the $i$-th solution. This function aims to minimize the discrepancy between the model’s output and the degree of correctness of each answer, across all questions and solution paths.
The output of this trained OSV, $f_{\bm{\theta}}(q;S^{(1:t)})$, approximates the expected value of $y_{i}$, i.e., the probability of correctness from the current step $t$ to the final answer. The following theorem [22] clarifies this claim:
@@
Proof of this theorem can be obtained by optimizing the MSE loss defined in the overall objective function (2), with further details provided in [22].
Moreover, we calculate the relative change in the confidence score between steps, represented as:
@@
$\Delta_{conf}^{t}$ represents the relative variation in the model’s confidence score from step $t$ to step $t+1$. A negative value of $\Delta_{conf}^{t}$ signifies a reduced confidence in achieving a correct answer after incorporating information from the $(t+1)$-th step.
We denote the process label for the $t$-th step as $y_{i}^{t}$ and $\theta$ as the variation threshold. Our automated process labelling AutoCV follows the “first error location” strategy [15] as following:
1. If $\Delta_{conf}^{t}>\theta$, then $y_{i}^{t}=1$;
2. Otherwise, $y_{i}^{t}=0$ and $\forall t^{\prime}>t$, $y_{i}^{t^{\prime}}=0$ regardless of $\Delta_{conf}^{t^{\prime}}$, i.e., if any step contains an error in a reasoning problem, the subsequent steps and the final result are considered incorrect.
Process-Supervision
We further fine-tune the OSV in a process-supervision manner.
$\mathcal{L}_{proc}(S_{i}^{(1:t)},y_{i}^{t};q)=\left(f_{\bm{\theta}}(S_{i}^{(1:t)};q)-y_{i}^{t}\right)^{2}.$ (4)
@@ 4 Preliminary Findings
In this section, we present our findings aiming to validate two key aspects: In Section 4.1, we present a comprehensive analysis of the OSV model, i.e., to validate that the initially trained OSV model is effective and robust. In Section 4.2, we further introduce a self-designed benchmark for process errors and calculate $\Delta_{conf}^{t}$ to detect these errors, i.e., to demonstrate the effectiveness and reliability of relative confidence variation in the proposed method. The validation of these two components serves as a foundation for automatic process labeling via AutoCV, as described in Section 3.3.
4.1 Experiment on Outcome Supervised Verifier Performance
@@
To evaluate the OSV, we measure its ability to select a sample with the correct final answer from samples generated by various LLMs, denoted as the Response Generator. Specifically, we assess the effectiveness of outcome supervision on two models: Phi2 (OSV-Phi[36] and Mistral-7B (OSV-Mistral[33]. To explore the scalability of this outcome-supervised verifier effect, we choose response generators of varying scales, ranging from 7B to 72B parameters, i.e., Mistral-7B-Instruct (Mistral-7B[33], Mixtral 8 $\times$ 7B (Mixtral-8$\times$7B[32] and Qwen-72B-Chat (Qwen-72B[35]. We employ the metric pass@$k$ [37] to assess the performance of those response generators, defined as the scenario where at least one correct instance is within the model’s first $k$ attempts. We evaluate the OSV’s generalized selection capability across different LLM scales.
Table 3: Performance of OSV models across different configurations. The selection of trained OSV models is based on 5 outputs from response generators.
Response Generator Pass@1 Pass@5 SC OSV-Mistral OSV-Phi
Mistral-7B 42.08 69.90 50.03 60.72 52.61
Mixtral-8$\times$7B 62.55 82.31 69.06 74.07 69.37
Qwen-72B 77.03 91.13 81.27 85.00 84.19
@@
The performance disparity among the verifiers can be primarily attributed to variations in model sizes and the accuracy of their training data. Specifically, the limited selection capability of the OSV-Phi model stems from discrepancies in the accuracy of its training data. Table Table 4 presents a consolidated view of the model sizes alongside the accuracy metrics of their outcome supervision training data.
@@
OSV-Mistral 7B 0.9914 100
OSV-Phi 2.7B 0.9605 100
In our content, we select OSV-Mistral as the OSV model among other experiment settings due to its superior performance, as demonstrated in Table ??????3.
4.2 Detecting Hallucination During Math Reasoning
In this section, we verify the effectiveness and reliability of our method AutoCV. Specifically, We calculate $\Delta_{conf}^{t}$ to identify inaccuracies in the process, as outlined in Eq. (3).
In Section 4.2.1, we introduce the concept of Process Calculation Hallucination (PCH) and establish a preliminary benchmark. In Section 4.2.2, we assess the performance of calculating $\Delta_{conf}^{t}$ to detect these PCH against our established benchmark.
4.2.1 Process Calculation Hallucination
We outline a method for identifying instances of PCH, which we define as occurrences where the numerical values on either side of an equals sign within a mathematical expression do not align. This misalignment indicates a breakdown in logical reasoning, categorizing the instance as a hallucination in the context of mathematical problem-solving. This process establishes a benchmark for PCH detection with more details in Appendix E.
Process Calculation Hallucination Detection
To identify hallucinations in mathematical reasoning, we monitor the relative confidence variations between the intermediate steps as defined in Eq. (3). If $\Delta_{conf}^{t}\leq\theta$, the corresponding step would be viewed as “incorrect”. Samples of detection are presented in Figure 2 of Appendix E.

### v4

# AutoPSV: Automated Process-Supervised Verifier

@@ Abstract
In this work, we propose a novel method named Automated Process-Supervised Verifier (AutoPSV) to enhance the reasoning capabilities of large language models (LLMs) by automatically annotating the reasoning steps. AutoPSV begins by training a verification model on the correctness of final answers, enabling it to generate automatic process annotations. This verification model assigns a confidence score to each reasoning step, indicating the probability of arriving at the correct final answer from that point onward. We detect relative changes in the verification’s confidence scores across reasoning steps to automatically annotate the reasoning process, enabling error detection even in scenarios where ground truth answers are unavailable. This alleviates the need for numerous manual annotations or the high computational costs associated with model-induced annotation approaches. We experimentally validate that the step-level confidence changes learned by the verification model trained on the final answer correctness can effectively identify errors in the reasoning steps. We demonstrate that the verification model, when trained on process annotations generated by AutoPSV, exhibits improved performance in selecting correct answers from multiple LLM-generated outputs. Notably, we achieve substantial improvements across five datasets in mathematics and commonsense reasoning. The source code of AutoPSV is available at https://github.com/rookie-joe/AutoPSV.
@@ 1 Introduction
Large language models (LLMs) have shown impressive performance on various reasoning tasks [1, 2, 3, 4]. Prior efforts primarily focus on specific prompting techniques, such as few-shot prompting with intermediate steps and augmented demonstrations [5, 6, 7, 8]. While these methods have shown promise, their effectiveness is often task-specific, and designing prompts can be labor-intensive, leading to inconsistent results [9, 10]. Another approach to improve reasoning in LLMs is through instruction tuning or knowledge distillation [11, 12, 13, 14]. These methods typically involve fine-tuning LLMs and require a large set of examples annotated with chain-of-thoughts (CoT; [5]). However, these approaches can be resource-intensive and may not always produce reliable results.
@@
To address these challenges, verification techniques have emerged as a promising solution [15, 16]. Verification models are trained to evaluate and potentially correct the reasoning process generated by LLMs. This approach aims to mitigate the risk of relying solely on the top-1 result, which may not always be reliable [17, 18]. By reranking candidate responses, verification models can ensure higher accuracy and consistency in LLM outputs, and provide valuable feedback for improving LLMs [19, 20] further.
@@
Verification models generally fall into two training paradigms: outcome supervision and process supervision. In outcome supervision, the training annotations rely on the correctness of the final answer [21, 22], while in process supervision, the annotations are based on evaluations of each reasoning step [23, 19]. However, process supervision is demanding in terms of annotations. Typically, it relies on either expensive and highly skilled human evaluators [23, 16] or model-induced process annotations [18, 17] to estimate the future correctness of the current reasoning step using Monte Carlo tree search [24, 25]. In contrast, outcome supervision only requires annotations for the output, making it more economical in terms of annotation effort but less effective. That being said when answers involve multiple reasoning paths, all aforementioned model-induced methods require numerous samples to ensure accurate estimations.
@@
In this paper, we introduce Automated Process-Supervised Verfier (AutoPSV), a novel approach that synergistically combines the strengths of both process supervision and output supervision. Our method begins by training an outcome-supervised verification model using outcome supervision annotations. This model then assigns confidence scores to each intermediate reasoning step, estimating their likelihood of contributing to a correct final answer. A distinguishing feature of AutoPSV is its ability to automatically generate process annotations through relative step-level confidence change analysis, significantly reducing annotation effort while maintaining supervision quality without requiring ground truth answers. These automatically generated process annotations subsequently serve as training data for developing an enhanced verification model that leverages both process-level and outcome-level supervision signals. The complete framework of AutoPSV is illustrated in Figure 1. We conduct extensive experiments across five datasets, including mathematical reasoning benchmarks and commonsense reasoning tasks. The results demonstrate that our method effectively improves the reasoning capability of the model with our highly efficient labeling scheme for process supervision. Our contribution is summarized as follows:
@@
  • We introduce AutoPSV to automate the labeling of process data to enhance LLMs’ reasoning capabilities. By combining the strengths of output and process supervision, AutoPSV effectively identifies variations in model confidence to annotate the correctness of intermediate reasoning steps, enabling efficient automatic labeling for process supervision.
@@
  • Comprehensive experiments demonstrate that AutoPSV significantly improves the performance and scalability of verification models in mathematical and commonsense reasoning tasks. This approach greatly reduces the need for manual intervention and extensive computational resources, making it a valuable tool for enhancing LLM capabilities.
  • AutoPSV’s versatility is evident in its applicability to both labeled and unlabeled dataset settings after completing the training process. This flexibility and generalizability highlight the method’s potential for widespread adoption in various LLM applications.
@@
Figure 1: An overview of AutoPSV. It utilizes an outcome-supervised verifier to automatically generate process annotations for each reasoning step by detecting its own confidence variations, without relying on ground truth annotations. AutoPSV efficiently produces annotations serving as process supervision during the LLM training, which sidesteps costly annotations.
@@ 2 Related Works
To enhance the reasoning capabilities of LLMs, prior research primarily focuses on specific prompting techniques [26]. Existing efforts include few-shot prompting with intermediate steps augmented demonstrations [5, 6, 7, 27] or zero-shot prompting with specific instructions [28, 29]. Although these methods have shown promising results, their effectiveness is often constrained by their task-specific nature and the labour-intensive process of designing prompts, leading to inconsistent outcomes across different tasks [9, 10]. Another strategy to facilitate reasoning involves instruction tuning or knowledge distillation, which elicits reasoning paths from LLMs without explicit prompting [11, 12, 13, 30]. These approaches typically involve resource-intensive fine-tuning over LLMs and require a large set of examples annotated with chain-of-thoughts (CoT). Unlike methods that directly modify parameters or prompts, AutoPSV focuses on training an additional verification model to select the desired output from the original model’s output. This approach is further discussed in the context of process supervision in the following paragraph.
@@
Recent efforts have focused on enhancing the reasoning capabilities of LLMs through the use of verifiers to select the best answer from multiple candidates. There are two main types of verifiers: the Outcome-Supervised Verifier (OSV) and the Process-Supervised Verifier (PSV). The OSV is supervised with a signal based on the final answer [21, 22], while the PSV is with detailed feedback which requires evaluating individual reasoning steps [15, 19, 16, 23]. Despite the time-consuming annotation cost, the PSV offers several advantages that make it preferable to the OSV. It can provide fine-grained feedback by pinpointing the location of errors, which is valuable for reinforcement learning and automatic correction [16, 20]. To alleviate the extensive human annotation, recent approaches [17, 18] propose a machine annotation framework using Monte Carlo Tree Search [24, 25]. This process demands a lot of computing resources, potentially imposing a limitation on the usage. AutoPSV is more efficient, as it utilizes an outcome-supervised verification model to assign confidence to each reasoning step and calculate relative step-level confidence changes, eliminating the need for additional sampling or manual labeling.
3 AutoPSV
In this section, we introduce the main problem that this paper focuses on in Section 3.1. We then discuss the motivation behind why we believe it is necessary to train a verification model in Section 3.2. Finally, we describe how we accomplish the transition from outcome supervision to process supervision during the training of the verification model in Section 3.3.
@@ 3.1 Problem Setting
Objective
Our research addresses the challenge of response selection from multiple candidates generated by a Large Language Model (LLM). Specifically, given an LLM acting as a response generator, we seek to develop an effective method for identifying the correct response among multiple generated solutions. Our primary goal is to maximize the probability of selecting an accurate solution from the available candidates.
Notation
We define $q$ as the input question presented to the model, and $S^{(1:t)}_{i}$ as the sequence of intermediate reasoning steps up to step $t$ for the $i$-th solution to question $q$. The binary correctness label for the $i$-th solution is denoted as $y_{i}$, and $\mathcal{Q}$ represents our training question dataset. This formalization enables us to systematically approach the response selection problem while maintaining mathematical rigor in our methodology.
@@ 3.2 Motivation
Response selection methods can be broadly divided into two categories: models specifically fine-tuned for the selection task and those that employ various prompting strategies.
In our exploration, we initially investigate whether existing open-source LLMs could serve as effective selection agents to evaluate model outputs and choose the correct response without fine-tuning. We choose Mixtral-Instruct [31] as the response generator, and its response results are listed Table 1. Our objective focuses on identifying the correct response from five candidate solutions.
To establish robust conclusions, we evaluate selector models ranging from 7B to over 70B parameters, applying various prompting strategies. The results, tested on the GSM8K test set [21], are presented in Table 2. Notably, even models exceeding 70 billion parameters demonstrate suboptimal selection performance when relying solely on prompting without fine-tuning.
Table 1: Performance of Mixtral-Instruct on GSM8K. All results are reported in accuracy (%).
@@
Table 2: Comparison of different selection methods across various model sizes for selecting a response from candidate responses generated by Mixtral-Instruct. All results are reported in accuracy (%).
@@
Based on these findings, our research focuses on training dedicated verification models and enhancing their response selection capabilities.
@@ 3.3 Training Methodology
AutoPSV integrates outcome supervision with process supervision to create an effective training methodology. Below, we detail each component of our approach.
Outcome-Supervision
We begin by training an outcome-supervised verification (OSV) model, denoted as $f_{\bm{\theta}}(\cdot)$, where $\bm{\theta}$ represents the optimized parameters. The training utilizes Mean Squared Error (MSE) loss for each solution step:
@@
The complete objective function across all training questions $\mathcal{Q}$ is:
@@
Where $n$ represents solutions per question and $m_{i}$ denotes steps in the $i$-th solution.
The OSV output approximates the expected correctness probability, as formalized in the following theorem:
@@
The proof follows from optimizing the MSE loss in equation (2), with details in [22].
Process-Supervision
We compute the relative confidence change between steps:
@@
$\Delta_{conf}^{t}$ represents the relative variation in the model’s confidence score from step $t$ to step $t+1$. A negative value of $\Delta_{conf}^{t}$ signifies a reduced confidence in achieving a correct answer after incorporating information from the $(t+1)$-th step. We denote the process label for the $t$-th step as $y_{i}^{t}$ and $\theta$ as the variation threshold.
For process labeling, we employ the "first error location" strategy [15] with threshold $\theta$:
  • If $\Delta_{conf}^{t}>\theta$: $y_{i}^{t}=1$
  • Otherwise: $y_{i}^{t}=0$ and $\forall t^{\prime}>t$, $y_{i}^{t^{\prime}}=0$
The process-supervision loss function is:
$\mathcal{L}_{proc}(S_{i}^{(1:t)},y_{i}^{t};q)=\left(f_{\bm{\theta}}(S_{i}^{(1:t)};q)-y_{i}^{t}\right)^{2}$ (4)
@@ 4 Preliminary Findings
In this section, we present our findings aiming to validate two key aspects: In Section 4.1, we present a comprehensive analysis of the OSV model, i.e., to validate that the initially trained OSV model is effective and robust. In Section 4.2, we further introduce a self-designed benchmark for process errors and calculate $\Delta_{conf}^{t}$ to detect these errors, i.e., to demonstrate the effectiveness and reliability of relative step-level confidence change in the proposed method. The validation of these two components serves as a foundation for automatic process labeling via AutoPSV, as described in Section 3.3.
4.1 Experiment on Outcome-Supervised Verifier Performance
@@
To evaluate the OSV, we measure its ability to select a sample with the correct final answer from samples generated by various LLMs, denoted as the Response Generator. Specifically, our task involves selecting the correct candidate from five responses. We assess the effectiveness of outcome supervision on two models: Phi2 (OSV (Phi)[35] and Mistral-7B (OSV (Mistral)[32]. To explore the scalability of this outcome-supervised verifier effect, we choose Response Generators of varying scales, ranging from 7B to 72B parameters, i.e., Mistral-7B-Instruct (Mistral-Instruct[32], Mixtral 8 $\times$ 7B (Mixtral-Instruct[31] and Qwen-72B-Chat (Qwen[34]. This allows us to check the OSV’s generalized ranking capability across different LLM scales.
Table 3: Performance of OSV models across different configurations.
Response Generator Pass@1 Pass@5 SC OSV (Mistral) OSV (Phi)
Mistral-Instruct 42.08 69.90 50.03 60.72 52.61
Mixtral-Instruct 62.55 82.31 69.06 74.07 69.37
Qwen 77.03 91.13 81.27 85.00 84.19
@@
The performance disparity among the verifiers can be attributed primarily to variations in model sizes and the quality of their training data. It’s important to note that the OSV model is continuously trained from the GSM8K fine-tuned model with the addition of a value head. This means that the training data for each OSV model is generated from its corresponding fine-tuned base model. Specifically, the training data for OSV (Mistral) is generated from the fine-tuned Mistral model, while the training data for OSV (Phi) is generated from the fine-tuned Phi model.
@@
OSV (Mistral) 7B 0.9914 100
OSV (Phi) 2.7B 0.9605 100
Table 4 presents a consolidated view of the model sizes along with the precision metrics of their outcome supervision training data.
It is worth noting that while both models are trained on the same quantity of data per question (100 samples), the quality of this data differs due to the capabilities of their respective base models. The Mistral model, being larger and potentially more capable, generates higher quality training data for its OSV, which in turn leads to better performance. In our content, we select the OSV (Mistral) model as the OSV model among other experiment settings due to its superior performance, as demonstrated Table 3.
4.2 Detecting Calculation Error During Math Reasoning
In this section, we verify the effectiveness and reliability of our method AutoPSV. Specifically, We calculate $\Delta_{conf}^{t}$ to identify inaccuracies in the process, as outlined in Eq. (3).
In Section 4.2.1, we introduce the concept of math calculation error and establish a preliminary benchmark. In Section 4.2.2, we assess the performance of calculating $\Delta_{conf}^{t}$ to detect calculation errors against our established benchmark.
4.2.1 Math Calculation Error
We outline a method for identifying instances of math calculation error, which we define as occurrences where the numerical values on either side of an equals sign within a mathematical expression do not align. This misalignment indicates a breakdown in logical reasoning, categorizing the instance as a calculation error in the context of mathematical problem-solving. This process establishes a benchmark for math calculation error detection with more details in Appendix E.
Math Calculation Error Detection
To identify calculation errors in mathematical reasoning, we monitor the relative step-level confidence changes between the intermediate steps as defined in Eq. (3). if $\Delta_{conf}^{t}\leq\theta$, we view the step as “incorrect”. We also provide a detection example in Figure 2 for better understanding.